Please help answer these questions

  1. A child with severe poison ivy weighs 20 kg and Benadryl PO 4 mg/kg/day is ordered every 6 hours. Benadryl is available as a 12 mg/5 ml solution. What dose should be given at 6:00 am? Give in mg and ml.
  2. The prescriber ordered Claforan (cefotaxime) 700 mg IM every 12 hours. The 1 gm vial of Claforan is in a powder form. The package insert states for IM injection, add 3.3 ml of sterile water for injection for an approximate volume of 3 mL containing 300 mg / ml. How many milliliters will you give the patients?
  3. Order: Lanoxin (Digoxin) elixir 0.25 mg PO every 12 hours. The child weighs 80 lbs and the recommended maintenance dose is 6–9 mcg / kg /day. a. What is the minimum daily maintenance dosage in mg / day? b. What is the maximum daily maintenance dosage in mg / day?
  4. Directions for reconstitution of mezlocillin of IV injection: "For IV use reconstitute with 30mL sterile water for injection. Each 30 mL of reconstituted solution will contain 3 grams of mezlocillin." The client is to receive 1.5 grams of mezlocillin IV every 6 hours. How many milliliters would the nurse administer to the client?

  5. Directions for reconstitution of ampicillin sodium for injection: "For IM use add 3.5 mL sterile water for injection. Resulting solution contains 250 mg ampicillin sodium per mL. The vial contains ampicillin sodium equivalent to 1 gram. Usual adult dosage of ampicillin is 250 mg–500 mg IM every 6 hours." How much diluent is added to the vial to prepare the medication for IM use?

  6. Directions for reconstitution of ampicillin sodium for injection: "For IM use add 3.5 mL sterile water for injection. Resulting solution contains 250 mg ampicillin sodium per mL. The vial contains ampicillin sodium equivalent to 1 gram. Usual adult dosage of ampicillin is 250 mg–500 mg IM every 6 hours." How much diluent is added to the vial to prepare the medication for IM use?

  7. A child weighs 18 kg and has an order for Zithromax 10 mg/kg po stat. The recommended dosage range on day one is 10 mg/kg. The oral suspension is dispensed as 200 mg in 5 mL. How many milliliters will you administer in one dose?

  8. Directions for reconstitution of oxacillin sodium for injection: "For IM use add 5.7 mL sterile water for injection. Resulting solution contains 250 mg oxacillin sodium per 1.5 mL. The vial contains oxacillin sodium equivalent to 1 gram. Usual adult dosage of oxacillin sodium is 250 mg–500 mg IM every 4 to 6 hours." The client is to receive 400 mg IM every 6 hours. How many milliliters would the nurse administer to the client?

  9. Order states administer 1000 ml D5NS IV from 4 am to 4 pm using a macro drip (10 gtts.ml). How many gtts/min?

  10. Order states diltiazem (Cardizem) 10 mg/hour IV. Supplied: diltiazem (Cardizem) 125 mg in 100 mL D5W and will infuse per infusion pump. How many mls/hour?

  11. Order states Tobramycin 50 mg IV every 8 hours. Reconstitute with 4.5 mls sterile water to yield 80 mg in 2 ml. Add to 100 mL D5W and infuse over 60 minutes. Use macrodrip tubing at 15 gtts/ml. What is the rate?
Answer #1

●Given

Order:Benadryl

Route:oral

Frequency:every six hourly (four time a day)

Weight:29kg

Dose:4mg/kg/day (that is 1kg=4mg/day)

When 1kg=4mg

Then 20kg=??mg

Cross multiply and divide

=20×4÷1

=20×4

=80mg

80mg/day

So one dose 80÷4=20mg

Available:12mg/5ml (that is 12mg=5ml)

So 80mg=??ml

Cross multiply and divide

=80×5÷12

=400÷12

=33.3 or 33ml

So one dose 33÷4=8.25ml

Ans:20mg and 8.25ml

●Given

Order:Claforan

Route:IN

Frequency :12 hourly (twice a day)

After dilution 3ml contains 300mg/ml

So when 300mg=1ml

Then 700mg=??ml

Cross multiply and divide

=700×1÷300

=700÷300

=2.3ml

Ans:2.3 ml to be administered

●Given

Order:Lanoxin

Dose:0.25mg

Frequency:twice a day

Route oral

Weight:80lbs (0.454×80=36.3 kg on rounding off it is 33kg)

Maintainace dose is 6_9mcg/kg/day

So 1kg=6mcg/day

Then 33kg=??mcg/day

Cross multiply and divide

=33×6÷1

=33×6

=198mcg/day

To convert to mg divide by 1000

=198÷1000

=0.198 mg/day

So if it is 9mcg/kg/day

Then 33kg=??mcg/day

Cross multiply and divide

=33×9÷1

=33×9

=297mcg/day

To convert into mg divide by 1000

=297÷1000

=0.297mg/day

Ans:a.The minimum daily maintenance dose is0.198mg/day

b.The maximum daily maintenance dose is 0.297mg/day

●Given

Order:Mezlocillin

Route:IV

Frequency:Four times a day

Available:30ml=3gm Mezlocillin

Dose:1.5mg

When 3mg=30ml

Then 1.5mg=??ml

Cross multiply and divide

=1.5×30÷3

=45÷3

=15

Ans:The nurse will administer 15ml